On January 1, 2021, prison granted a P20 million loan to Break Company. The loan is repayable by the on equal annual instalments of P4.80 million over a five-year term. The effective interest rate that Prison charges the borrower is 6.4% per annum comprising 4% risk-free rate and 2.4% for credit risk. Prison estimate that there is a 70% chance that the loan will not default, a 20% chance that the loand default and the expected cash flow in each year us P3.60 million; and a 10% chance that the loan defaults and the expected cash flow in each year is P3.00 million.

Required:

1. Compute for the amount of cash shortfall for the five year period.

2. Compute for the probability weighted cash shortfall

3. Compute for the lifetime expected credit loss

4. Compute for the 12-month expected credit loss

5. Prepare the journal entry in 2021.



1
Expert's answer
2021-06-07T06:57:02-0400

Solution:

1.). Amount of cash shortfall:

Expected total annual instalments for the five years = 4.8m "\times" 5 = 24m

Cash flow expected using effective interest rate = "20m + ((20m\times 6.4\%)\times5) = 20m + 6.4m = 26.4m"

Amount of cash shortfall = 26.4m – 24m = 2.4m


2.). Probability weighted cash shortfall:

First derive the weighted probability:

Probability 1 = 70% = 0.7

Probability 2 = 20% = 0.2

Probability 3 = 10% = 0.1

Total weights          = 1

Weighted probability = Total weights / outcome = "\frac{1}{3}"


Probability weighted cash shortfall = "\frac{1}{3} \times 24m = 0.8m"


3.). The lifetime expected credit loss:

Expected Credit Loss (ECL) = The probability weighted estimate of all credit losses (the present value of all cash shortfalls over the expected life of the financial instrument.

ECL = Exposure At Default"\times" Probability

Exposure At Default 1 = 20m

Exposure At Default 2 = 3.5m"\times" 5 = 18m

Exposure At Default 3 = 3m"\times" 5 = 15m

ECL = (20m"\times" 0.3) + (18m"\times" 0.2) + (15m"\times" 0.1)

ECL = 6m + 3.6m + 1.5m = 11.1m

The lifetime expected credit loss = 11.1m


4.). 12 month expected credit loss:

Expected Credit Loss (ECL) = The probability weighted estimate of all credit losses (the present value of all cash shortfalls over the expected life of the financial instrument.

ECL = Exposure at Default"\times" Probability

Exposure At Default 1 = 4.8m

Exposure At Default 2 = 3.5m

Exposure At Default 3 = 3m

ECL = (4.8m"\times" 0.3) + (3.5m"\times" 0.2) + (3m"\times" 0.1)

ECL = 1.44m + 0.7m + 0.3m = 2.44m

12 month expected credit loss = 2.44m


5.). The journal entry for 2021 will be as follows:

Dr. Impairment loss in profit and loss 2.44m

Cr. Loss allowance in financial position 2.44m

(To recognize 12-moth expected credit loss)